OpenStudy (linn99123):

Factor the trinomial: 2x2 + 7x + 3 can someone teach me how to do it?

hartnn (hartnn):

Hi :) Can you find 2 numbers whose sum is +7 and product is \(2 \times 3= 6\) ?

Elsa213 (elsa213):

cx good job hartnn c:

OpenStudy (linn99123):

im trying to figure it out so do i put 2 times 2 to get 6

OpenStudy (linn99123):

im sorry im so confused

OpenStudy (linn99123):

nvm ill do this later im not feeling so well thanks anyways

Elsa213 (elsa213):

2*2=4

OpenStudy (linn99123):

yea ik i just made a fool of my self elsa /.\

hartnn (hartnn):

In general, for \(ax^2+bx+c\), we find 2 numbers whose sum is +b and product is ac. and 2 times 2 isn't 6 :) What are the factors of 6 ?

Elsa213 (elsa213):

(~;-;)~ dun worry i do dat all du time

Elsa213 (elsa213):

3*2=6 c:

OpenStudy (linn99123):

i already left the test.. thanks anyways @hartnn

OpenStudy (linn99123):

im dealing with some stuff so ill be back later.

Elsa213 (elsa213):

oooo like wut? maybe i can help :D

hartnn (hartnn):

Take care linn! let me know when you wanna learn :)

OpenStudy (linn99123):

nah its just personal problems i need to get over

OpenStudy (linn99123):

im about to log ill ttyl elsa

Elsa213 (elsa213):

oki bye linn c: